Problem 3. After plotting demand for four periods, an emergency room manager has concluded that a trend- adjusted exponential smoothing model is appropriate to predict future demand. The initial estimate of trend is based on the net change of 30 for the three periods from 1 to 4, for an average of +10 units. Use α-Sand β,4, and TAF of 250 for period 5. Obtain forecasts for periods 6 through 10. Period Actual 210 224 229 240 255 Period Actual 265 272 285 10 Problem 4 A producer of felt-tip pens from its marketing department. Fixed costs of tion, and variable costs are 37 cents per pen. a. Find the break-even quantity if pens sell for $1 each. b. At what price must pens be sold to obtain a monthly proft of $15,000, assuming has received a forecast of demand of 30,000 pens for the coming month $25,000 per month are allocated to the felt-tip opera that estimated FS F7 F8
Answer #1

4a. Break even quantity = Fixed costs/(sale price per unit - variable cost per unit)

= $25,000/($1 per pen - $0.37 per pen)

= $25,000/$0.63 per pen

= 39,682.54 pens. (This can be rounded off to 39,683 pens)

b. Now estimated demand = 30,000 and monthly profit required = $15,000. Let the price be $x per pen

Hence total revenue = 30,000x and total cost = 25000+(0.37*30000)

Profit = revenue - costs

= 30,000x - 25000-(0.37*30000)

Thus 30,000x - 25000-(0.37*30000) = 15000

or 30,000x = 51,100

or x = 51,100/30,000

$1.70 per pen